# LLM4SBR: A Lightweight and Effective Framework for Integrating Large Language Models in Session-based Recommendation

### v1

@@ 4.5. Space Occupation Experiment and Analysis
SR-GNN 1,314 1,224
LLM4SBR(SR-GNN) 1,330 1,304
Recommendation models based on LLM often require a large amount of video memory. To explore the spatial effectiveness of the LLM4SBR framework, we recorded the memory usage of SR-GNN and LLM4SBR (SR-GNN).
The results are shown in Table 3, the memory usage rates of the two are very close. Through the two-stage strategy, LLM inference is performed only once in the first stage, and then the inference results are stored in an external file. During the second stage of SBR model training, the memory consumption is limited to the original SBR model, the ID sequence data, and the pre-stored inference embeddings, significantly reducing memory usage. Additionally, since the second stage only requires parsing the stored inference embeddings into tensors, its increased time complexity is $O(1)$, the model’s time complexity mainly depends on the original SBR model. Taking into account time and space factors, the LLM4SBR framework we proposed can be implemented in industrial environments.
@@ 5. Conclusions and future work
In this paper, we explore the feasibility of combining LLM with SBR models while considering both effectiveness and efficiency. In short sequence data, LLM can infer preferences directly leveraging its language understanding capability without fine-tuning. This approach is more efficient in utilizing information compared to encoding text data into embeddings for training, and it allows us to place LLM and SBR in separate stages, greatly reducing training costs. Regarding the LLM hallucination, we found that it can be corrected through the similarity of text embeddings, and enhancement with similar samples can improve the diversity of inference results to a certain extent.
In addition, we propose a scalable two-stage LLM enhancement framework (LLM4SBR) tailored for SBR. Specifically, in the semantic reasoning phase, we utilize LLM as the inference engine, designing prompt-guided inference processes from different perspectives and leveraging an intent localization module to eliminate LLM hallucinations and achieve semantic enhancement. In the representation enhancement stage, we perform fine-grained alignment and uniformity of text embeddings and session embeddings from different perspectives. This effectively facilitates the fusion of representations from different modalities, thereby enhancing the final session representation. Extensive experiments have demonstrated the effectiveness of the LLM4SBR framework, which significantly enhances most SBR models while also improving model interpretability and enhancing the diversity of candidate selection.
For future work, we will continue exploring whether adding additional LLM inference perspectives can yield greater benefits, as well as assessing the effectiveness of utilizing LLM Agent for logical reasoning. In addition, we also want to explore the application of other downstream tasks combined with LLM. Finally, we hope for this work to open up new avenues in SBR research, accelerating deeper exploration into the integration of LLM with RS.

### v2

@@ 5.5. Model training space occupancy experiment and Time complexity analysis (RQ4)
Model SR-GNN LLM4SBR (SR-GNN) SR-GNN LLM4SBR (SR-GNN)
GPU Usage (MB) 1,262 1,282 1,290 1,324
Single Epoch Time (s) 398.08 1490.49 124.89 350.80
Training Time (s) 2085.67 2987.54 932.10 1406.79
Training LLM-based recommendation models typically require a significant amount of GPU resources and longer training time. To explore the spatial and temporal effectiveness of the LLM4SBR framework, We record the GPU usage during training, the training time for a single epoch, and the training time required to achieve the best performance for SR-GNN and LLM4SBR (SR-GNN).
The results are shown in Table 5. First, we can observe that the GPU occupancy rates of the original SR-GNN and the model using the LLM4SBR framework are very close. We believe that this is because the LLM4SBR framework only requires model training in the second stage, and this process does not require the participation of LLM. The SBR model only needs to parse the LLM inference results into tensor form to participate in training, thus saving a lot of training time and GPU resources. Second, looking at the combined single epoch training time and the time for the model training to reach optimal performance, it can be observed that although the training time of the model using the LLMSBR framework for a single epoch maybe 3-4 times longer than that of the original model, the training time to reach the optimal performance is not much different in comparison with the original model, which illustrates the fact that although the model using the LLM4SBR framework will increase the training time for a single epoch, the model can converge faster and reach optimal performance.
In addition, we analyze the time complexity of the model. In the first stage of LLM4SBR, each prompt corresponds to a conversation sequence. It is assumed that there are $n$ such prompts, and the length of each prompt is quite short compared to n, which means that the time for a single inference can be regarded as constant time $O(1)$. Given that there are $n$ prompts, each of which requires two inferences, the total time complexity of the entire stage is $O(2\times n\times 1)=O(2n)=O(n)$. In the second stage of SBR model training, since the second stage only needs to parse the stored inference embedding into tensors, its added time complexity is $O(n)$. The time complexity of this phase is $O(SBR+n)$, and $O(SBR)$ usually lies between $[n,n^{2}]$, so the time complexity of the LLM4SBR mainly depends on the original SBR model. Therefore, the total time complexity of the LLMSBR framework is $O(LLM4SBR)=O(SBR+n)\approx O(SBR)$, which depends on the time complexity of SBR and the size of the dataset.
In summary, the added space and time costs are small compared to the huge performance gains of the LLM4SBR framework.
Figure 8. The inference results of LLM with different parameter quantities (7B, 32B, 72B, 110B) for short-term and long-term views.
5.6. Case Study
In this section, we analyze a specific case and examine the level of understanding and inference results of prompts by LLMs with different parameter quantities.
As shown in Figure 8, we generated prompts from both short-term and long-term views for the text of the same session sequence and used LLMs ( 7B, 32B, 72B, and 110B ) for inference. The results are shown on the right of the figure, where ”S” represents the inference result of short-term interest and ”L” represents the inference result of long-term interest. We did not remove the text explanations generated by LLMs here. We summarize the observed phenomena and analysis results as follows:
  • All LLMs can respond to different view prompts, even those with the smallest parameter size. In Figure 8, we highlight in red the view-limiting qualifier in the prompt and the fields that respond to the view in the reply. We can see that all four LLMs in the figure understand the view-limiting qualifier in the prompt and make corresponding inferences in the response. In the inference results of the same LLM, the inference results of long-term interests and short-term interests are not the same. These phenomena indirectly prove that the inferencing ability of current general-purpose LLMs has been greatly improved and that they can infer reasonably within the original knowledge even without the fine-tuning of domain-specific knowledge.
  • The number of parameters in an LLM affects the inference ability and scope of a generalized LLM. In the figure, we use orange to indicate that the inference item exists in the user sequence, and blue to indicate that the inference item does not exist in the user sequence. We can observe that LLMs of 7B tend to infer item representatives of users’ long-term/short-term interests from the current sequence item set. In contrast, larger LLMs may generate inference results beyond the scope due to having more built-in knowledge. However, the LLM4SBR framework we proposed has already considered this situation and designed an intent localization module that can effectively constrain this LLM hallucination phenomenon. Furthermore, as shown in Table 3, even the 7B model can still bring significant performance improvements to the SBR model. The LLM4SBR framework allows free substitution of LLMs so that the appropriate LLM can be selected for inference depending on the actual requirements or computational resources, which provides a very high degree of flexibility and compatibility.
@@ 6. Conclusions and Future work
In this paper, we propose a scalable two-stage LLM enhancement framework (LLM4SBR) tailored for SBR. This approach is more efficient in utilizing information compared to encoding text data into embeddings for training, and it allows us to place LLM and SBR in separate stages, greatly reducing training costs. Specifically, in the semantic inference phase, we utilize LLM as the inference engine, designing prompt-guided inference processes from different views and leveraging an intent localization module to alleviate LLM hallucinations and enhance semantic. In the representation enhancement stage, we perform fine-grained alignment and uniformity of text embeddings and session embeddings from different views. This effectively facilitates the fusion of representations from different modalities, thereby enhancing the final session representation. Extensive experiments have demonstrated the effectiveness of the LLM4SBR framework, which significantly enhances most SBR models while also improving model interpretability and enhancing the diversity of candidate selection.
For future work, we will continue exploring whether adding additional LLM inference views can yield greater benefits, as well as assessing the effectiveness of utilizing LLM Agent for logical inference. In addition, we also want to explore the application of other downstream tasks combined with LLM. Finally, we hope for this work to open up new avenues in SBR research, accelerating deeper exploration into the integration of LLM with RS.
Acknowledgements.
This work is supported by the National Key Research and Development Program of China under 2022YFB3104702. This work is also supported by the National Natural Science Foundation of China under grants 62272262, 72442026, 72342032, 72074036, and 62072060.